# Learnability Window in Gated Recurrent Neural Networks

### 20 Mar 2026 v7

@@ 5.4 Finite-sample analysis
We derive the finite-sample requirements for detecting a lag-$\ell$ dependency from noisy gradient information.
@@ 5.4.1 Statistical model
[…]
We extend the zero-mean formulation of [49] to a two-parameter location family, allowing small shifts in the mean corresponding to the presence or absence of a detectable signal. At lag $\ell$, the averaged matched statistic $\widehat{S}_{N}(\ell)$ over $N$ independent training sequences is modeled as
$\widehat{S}_{N}(\ell)\;\big|\;B\;\sim\;\mathcal{S}\alpha\mathcal{S}\!\Big(\theta_{\text{out}},\;\sigma_{\alpha}(\ell)/N^{1-1/\alpha}\Big),\qquad\theta_{\text{out}}\!\in\!\left\{+\tfrac{1}{2}\Delta(\ell),\;-\tfrac{1}{2}\Delta(\ell)\right\},$ (16)
where $\theta_{\text{out}}$ is the location parameter, the mean separation $\Delta(\ell)=\overline{m}_{\mu}(\ell)f(\ell)$ quantifies the strength of the lag-$\ell$ signal, $\sigma_{\alpha}(\ell)$ acts as a noise-scale proxy, and $B\in\{0,1\}$ describes the non-detection/detection outcome.
@@ 5.4.2 Information-theoretic bounds
[…]
$P_{\text{det}}=\mathcal{S}\alpha\mathcal{S}\!\Big({+}\tfrac{1}{2}\Delta(\ell),\,\sigma_{\alpha}(\ell)/N^{\,1-1/\alpha}\Big),\qquad P_{\text{non}}=\mathcal{S}\alpha\mathcal{S}\!\Big({-}\tfrac{1}{2}\Delta(\ell),\,\sigma_{\alpha}(\ell)/N^{\,1-1/\alpha}\Big),$
where $\Delta(\ell)=\overline{m}_{\mu}(\ell)\,f(\ell)$ denotes the mean separation.
For $\alpha>1$, symmetric $\alpha$–stable location families satisfy local asymptotic normality (LAN) under suitable normalization, enabling the use of information-theoretic bounds [23, 34, 42]. In the present setting, the empirical statistic $\widehat{S}_{N}(\ell)$ defines a triangular array of location models [52] whose scale shrinks with the number of sequences as $s_{N}=\sigma_{\alpha}(\ell)N^{1/\alpha-1}$. For such families, LAN holds with normalization rate $r_{N}=1/s_{N}$, which governs the detectability thresholds derived below. As a consequence, the log-likelihood ratio between $P_{\text{det}}$ and $P_{\text{non}}$ admits a quadratic expansion for small location shifts, yielding the lower bound for the Kullback-Leibler (KL) divergence
@@
This expression links sample complexity to the noise scale $\sigma_{\alpha}(\ell)$, the average alignment $\overline{m}_{\mu}(\ell)$, and the envelope $f(\ell)$. In contrast to the Gaussian case, the exponent $\alpha/(\alpha-1)>2$ for $\alpha<2$, reflecting the slower statistical concentration induced by heavy-tailed gradient noise. Full derivations of Eqs. (18) and (19) are given in Appendix G.
@@ 5.5 Learnability window and scaling laws
Throughout this section we adopt the corrected concentration exponent
$\kappa_{\alpha}\;=\;\frac{\alpha}{\alpha-1},\qquad\alpha>1,$ (20)
which governs the finite-sample detectability of lagged dependencies under $\alpha$-stable gradient noise when statistics are averaged over independent sequences. Equivalently, the noise floor of the empirical matched statistic decays as $N^{-1/\kappa_{\alpha}}=N^{-(1-1/\alpha)}$.
@@
In these expressions, the symbol $\asymp$ denotes equality at the level of asymptotic order. These scaling laws characterize how gating geometry and heavy-tailed noise jointly determine the large-lag behavior of temporal learnability. Exponential decay corresponds to rapid forgetting and only logarithmic growth of the horizon; polynomial decay yields algebraic horizon growth; and logarithmic decay represents a boundary regime with rapidly expanding $\mathcal{H}_{N}$. In all cases, smaller $\alpha$ compresses $\mathcal{H}_{N}$ by increasing $\kappa_{\alpha}$ and slowing statistical concentration. Full derivations and constant factors are provided in Appendix I.
@@ 5.6 Theoretical implications
The concentration exponent $\kappa_{\alpha}=\alpha/(\alpha-1)$ quantifies how rapidly noise averages out when statistics are aggregated across independent sequences. For Gaussian fluctuations ($\alpha=2$), one recovers the familiar concentration rate $N^{-1/2}$; for $\alpha<2$, the slower rate $N^{-1/\kappa_{\alpha}}=N^{-(1-1/\alpha)}$ substantially weakens the effective signal-to-noise ratio. Thus, even for identical gating geometry and identical envelopes $f(\ell)$, smaller values of $\alpha$ increase $\kappa_{\alpha}$, compress the learnability window, and raise the number of samples required for reliable detection.
@@
Architectures that permit slowly decaying envelopes are intrinsically capable of supporting long-range learnability under heavy-tailed noise. At the same time, such behavior is not guaranteed by architecture alone: training dynamics and noise statistics determine which decay regime is effectively realized. Conversely, training methods that reduce the impact of heavy-tailed fluctuations, such as clipping, normalization, or stabilizing adaptive optimizers, effectively increase $\alpha$, thereby reducing $\kappa_{\alpha}$ and improving statistical efficiency without altering model capacity.
@@
The learnability window $\mathcal{H}_{N}$ may vanish identically for certain architectures, meaning that $\mathcal{H}_{N}=0$ for all finite $N$. In this case, the envelope $f(\ell)$ falls below the statistical detectability threshold for every $\ell\geq 1$, and the corresponding sample complexity $N(\ell)$ diverges. This does not reflect an optimization failure or insufficient data, but an information-theoretic obstruction: the transported gradients carry no statistically usable signal at nonzero lags. Equivalently, the KL divergence between the detection and non-detection models vanishes, so no estimator can reliably distinguish the two hypotheses regardless of sample size.
@@ 6 Empirical validation
Accordingly, the experiments are designed to test the structural predictions of the theory. In particular, we aim to: (i) measure the empirical decay profiles of the envelope $f(\ell)$ and assess how they align with the canonical decay regimes considered by the theory; (ii) quantify how the resulting empirical learnability windows $\widehat{\mathcal{H}}_{N}$ vary with gating structure, optimizer dynamics, and sample size, and examine whether their scaling behavior is consistent with the regime implied by the observed envelope geometry; and (iii) investigate how the neuronwise time-scale spectra $\{\tau_{q}\}_{q}$ and the estimated gradient-noise statistics jointly relate to the empirically realized learnability regimes.
@@ 6.1 Experimental setup
After training, parameters are frozen and all diagnostics are computed on an independent set of $8000$ fresh sequences generated from the same data-generating process, with no overlap in sequences or random seeds. For each trained model, the transport factors $\Gamma^{(q)}_{t,\ell}$ are computed from the frozen gate activations using the architecture-specific expressions derived in Appendix B. For AdamW, the adaptive base rates $\Lambda^{(q)}_{r,\ell}$ are obtained from the Rayleigh projection (7); for SGD they reduce to the constant $\Lambda^{(q)}_{r,\ell}=\mu$. For each diagnostic lag $\ell$ and unit $q$, the magnitudes $|\mu_{t,\ell}^{(q)}|$ are averaged over time steps and independent sequences. The empirical envelope is then defined as
$\hat{f}(\ell)=\sum_{q}\Big\langle|\mu_{t,\ell}^{(q)}|\Big\rangle,$
@@
Empirical learnability window.
@@
where the right-hand side is a positive constant depending only on $\alpha$ and the error probability $\epsilon$. This reformulation makes explicit that the Fano bound constrains a single scalar quantity, i.e. the SNR, and that the (aggregated) detection threshold $\varepsilon$ affects only the absolute position of $\mathcal{H}_{N}$, not how it scales with $N$.
In practice, we estimate the SNR by substituting the empirical signal $|\widehat{\overline{m}}_{\mu}(\ell)|$, the estimated tail index $\hat{\alpha}$, and the estimated scale $\hat{\sigma}_{\alpha}(\ell)$:
$\widehat{\mathrm{SNR}}(\ell,N)\;=\;\frac{|\widehat{\overline{m}}_{\mu}(\ell)|\;N^{1-1/\hat{\alpha}}}{\hat{\sigma}_{\alpha}(\ell)}.$
Since $c_{\alpha}$ is an existence constant from the LAN expansion (Appendix F) and is not available in closed form, we set the detection threshold $\varepsilon$ directly. This preserves the full scaling structure of the theoretical bound (the concentration rate $N^{1-1/\alpha}$, the dependence on signal amplitude and noise scale) while replacing the single unknown constant with a user-specified parameter. The empirical learnability window is then
@@ 7 Discussion and future directions
This work develops a theoretical framework linking gating geometry, optimizer adaptation, heavy-tailed gradient noise, and the finite-horizon learnability of recurrent neural networks. The central message is that the generalized effective learning rates, rather than Jacobian stability in isolation, determine how much gradient signal remains statistically detectable across temporal lags. The decay profile of the envelope, together with the tail index $\alpha$, governs both sample-complexity requirements and the learnability window.
@@
The scaling laws make precise the statistical cost of each decay regime: exponential envelope decay entails exponential sample complexity, while polynomial decay entails only polynomial sample complexity. For any finite training budget, exponential forgetting rapidly exhausts the available statistical resources, whereas polynomial forgetting keeps long-range dependencies within reach. Heavy-tailed gradient fluctuations exacerbate this disparity by increasing $\kappa_{\alpha}$ and slowing statistical concentration. Heavy-tailed noise therefore acts as a viability constraint, disfavoring rapid forgetting and effectively selecting, among the dynamical regimes accessible to an architecture–optimizer pair, those that preserve signal over longer temporal scales.
We conjecture that this viability constraint induces a form of constrained self-organization: during training, recurrent networks—and potentially deep learning models more broadly—tend to evolve toward regimes of slower envelope decay in order to maintain learnability under their own gradient-noise statistics. This does not require explicit regularization; it follows from the fact that configurations with exponential forgetting become statistically unlearnable at long lags, creating implicit selection pressure toward slower decay. The achievable regime depends on the architecture–optimizer pair: constrained gating limits the accessible spectra, while flexible gating combined with adaptive optimization permits broader time-scale mixtures. This conjecture is consistent with our empirical observation that slow-decay regimes systematically co-occur with heavier-tailed gradient noise, whereas synchronized exponential regimes operate closer to Gaussian concentration.
@@
These observations highlight three broader implications. First, the temporal horizon of learnability is primarily controlled by the joint action of the envelope decay and optimizer adaptation. Second, the architecture–optimizer pair shapes the space of attainable time-scale spectra, biasing—but not guaranteeing—the emergence of heterogeneous dynamics that support long-range credit assignment. Third, under heavy-tailed gradient noise, increasing the number of training sequences yields diminishing returns for configurations operating in the exponential regime, whereas modifying the architecture–optimizer pair can alter the scaling class of temporal learnability entirely.

### 28 Mar 2026 v8

@@ 5.4 Finite-sample analysis
Here, we derive the finite-sample requirements for detecting a lag-$\ell$ dependency from noisy gradient information.
@@ 5.4.1 Statistical model
Empirical studies of SGD indicate that gradient fluctuations in deep networks are well-described by a $\mathcal{S}\alpha\mathcal{S}$ distribution with $1\leq\alpha\leq 2$, rather than a Gaussian one [48, 8]. $\mathcal{S}\alpha\mathcal{S}$ random variables form a family of heavy-tailed laws indexed by the tail index $\alpha\!\in\!(0,2]$. They are characterized by the characteristic function $\phi_{X}(t)=\exp\!\big(-\sigma^{\alpha}|t|^{\alpha}\big)$, where $\sigma>0$ is a scale parameter controlling dispersion. Their probability densities generally lack closed form except in special cases (Gaussian for $\alpha=2$ and Cauchy for $\alpha=1$), and decay with power-law tails as $p(x)\sim|x|^{-(1+\alpha)}$. For $\alpha<2$, these distributions have infinite variance; moreover, the first moment is finite only when $\alpha>1$.
Motivated by this literature, and noting that the matched statistic aggregates gradient-derived contributions through linear operations, we model its fluctuations with a $\mathcal{S}\alpha\mathcal{S}$ location family (see Appendix E for a detailed justification). At lag $\ell$, the finite-sample matched statistic $\widehat{S}_{N}(\ell)$ defined in Eq. (17) is modeled as
$\widehat{S}_{N}(\ell)\;\sim\;\mathcal{S}\alpha\mathcal{S}\!\Big(\theta_{\text{out}},\;\sigma_{\alpha}(\ell)/N^{1-1/\alpha}\Big),\qquad\theta_{\text{out}}\!\in\!\left\{+\tfrac{1}{2}\Delta(\ell),\;-\tfrac{1}{2}\Delta(\ell)\right\},$ (18)
where $\theta_{\text{out}}$ is the location parameter, the mean separation $\Delta(\ell)=\overline{m}_{\mu}(\ell)f(\ell)$ quantifies the strength of the lag-$\ell$ signal, and $\sigma_{\alpha}(\ell)$ acts as a noise-scale proxy.
@@ 5.4.2 Information-theoretic bounds
The learnability problem at lag $\ell$ is equivalent to distinguishing between the two location models in Eq. (18). We denote by $P_{\text{det}}$ and $P_{\text{non}}$ the cases $\theta_{\text{out}}=+\tfrac{1}{2}\Delta(\ell)$ and $\theta_{\text{out}}=-\tfrac{1}{2}\Delta(\ell)$, respectively.
For $\alpha>1$, symmetric $\alpha$–stable location families satisfy local asymptotic normality (LAN) under suitable normalization, enabling the use of information-theoretic bounds [23, 33, 42]. In the present setting, the finite-sample statistic $\widehat{S}_{N}(\ell)$ defines a triangular array of location models [51] whose scale shrinks with the number of sequences as $s_{N}=\sigma_{\alpha}(\ell)N^{1/\alpha-1}$. For such families, LAN holds with normalization rate $r_{N}=1/s_{N}$, which governs the detectability thresholds derived below. As a consequence, the log-likelihood ratio between $P_{\text{det}}$ and $P_{\text{non}}$ admits a quadratic expansion for small location shifts, yielding the lower bound for the Kullback-Leibler (KL) divergence
@@
This expression links sample complexity to the noise scale $\sigma_{\alpha}(\ell)$, the average alignment $\overline{m}_{\mu}(\ell)$, and the envelope $f(\ell)$. In contrast to the Gaussian case, the exponent $\alpha/(\alpha-1)>2$ for $\alpha<2$, reflecting the slower statistical concentration induced by heavy-tailed fluctuations. Full derivations of Eqs. (20) and (21) are given in Appendix G.
@@ 5.5 Learnability window and scaling laws
Throughout this section, we adopt the corrected concentration exponent $\kappa_{\alpha}=\alpha/(\alpha-1)$, $\alpha>1$, which governs the finite-sample detectability of lagged dependencies under $\alpha$-stable fluctuations when statistics are averaged over independent sequences. In this regime, the noise floor of the averaged statistic scales as $N^{-1/\kappa_{\alpha}}=N^{-(1-1/\alpha)}$.
@@
In these expressions, the symbol $\asymp$ denotes equality at the level of asymptotic order. [^4] We write $f\asymp g$ to mean that $c_{1}\,g\leq f\leq c_{2}\,g$ for positive constants $c_{1},c_{2}$ independent of the asymptotic variable. Unlike $f\sim g$ (which requires $f/g\to 1$), the symbol $\asymp$ preserves the correct scaling exponent while absorbing the architecture- and threshold-dependent multiplicative constants ($c_{\sigma},C_{\sigma},c_{m},C_{m},\kappa_{\alpha,\epsilon}$) that appear in the sandwich bounds of Appendix I. These scaling laws characterize how gating geometry and heavy-tailed noise jointly determine the large-lag behavior of temporal learnability. Exponential decay corresponds to rapid forgetting and only logarithmic growth of the horizon; polynomial decay yields algebraic horizon growth; and logarithmic decay represents a boundary regime with rapidly expanding $\mathcal{H}_{N}$. In all cases, smaller $\alpha$ compresses $\mathcal{H}_{N}$ by increasing $\kappa_{\alpha}$ and slowing statistical concentration. Full derivations and constant factors are provided in Appendix I.
@@ 5.6 Theoretical implications
The concentration exponent $\kappa_{\alpha}=\alpha/(\alpha-1)$ quantifies how rapidly noise averages out when statistics are aggregated across independent sequences. For Gaussian fluctuations ($\alpha=2$), one recovers the familiar concentration rate $N^{-1/2}$; for $\alpha<2$, the slower rate $N^{-1/\kappa_{\alpha}}=N^{-(1-1/\alpha)}$ substantially weakens the effective signal-to-noise ratio. Thus, holding the gating geometry, the optimizer-induced envelope $f(\ell)$, and the noise scale $\sigma_{\alpha}(\ell)$ fixed, smaller values of $\alpha$ increase $\kappa_{\alpha}$, compress the learnability window, and raise the number of samples required for reliable detection.
@@
Architectures that permit slowly decaying envelopes are, in principle, better positioned to support long-range learnability under adverse noise statistics. However, this behavior is not determined by architecture alone: training dynamics and the realized noise regime jointly determine which decay profile is effectively stabilized in practice. Accordingly, optimization choices that alter fluctuation structure or stabilize training can improve learnability either by modifying the realized envelope, by changing the effective noise scale, or by shifting the effective tail behavior of the aggregated statistic.
@@
In the limiting case where the effective lag signal vanishes identically at all nonzero lags, the learnability window collapses: $\mathcal{H}_{N}=0$ for all finite $N$. Equivalently, the KL divergence between the detection and non-detection models is zero, so no estimator can reliably distinguish the two hypotheses regardless of sample size. This is not an optimization failure or a data limitation, but an information-theoretic obstruction: the transported gradients carry no statistically usable signal at nonzero lags.
@@ 6 Empirical validation
Accordingly, the experiments are designed to test the structural predictions of the theory. In particular, we aim to: (i) measure the empirical decay profiles of the envelope $f(\ell)$ and assess how they align with the canonical decay regimes considered by the theory; (ii) quantify how the resulting empirical learnability windows $\widehat{\mathcal{H}}_{N}$ vary with gating structure, optimizer dynamics, and sample size, and examine whether their scaling behavior is consistent with the regime implied by the observed envelope decay; and (iii) investigate how the neuronwise time-scale spectra $\{\tau_{q}\}_{q}$ and the estimated matched statistic fluctuations jointly relate to the empirically realized learnability regimes.
@@ 6.1 Experimental setup
After training, parameters are frozen and all diagnostics are computed on an independent set of $8000$ fresh sequences generated from the same data-generating process, with no overlap in sequences. This isolates the diagnostic measurements from the optimization trajectory itself and probes the transport and noise structure induced by the trained model. For each trained model, the transport factors $\Gamma^{(q)}_{t,\ell}$ are computed from the frozen gate activations using the architecture-specific expressions derived in Appendix B. For AdamW, the adaptive base rates $\Lambda^{(q)}_{r,\ell}$ are obtained from the Rayleigh projection (7); for SGD they reduce to the constant $\Lambda^{(q)}_{r,\ell}=\mu$.
Empirical envelope.
For each diagnostic lag $\ell$ and unit $q$, we summarize the effective learning-rate magnitude by averaging $|\mu_{t,\ell}^{(q)}|$ over admissible time positions and independent diagnostic sequences. Writing $\langle\cdot\rangle_{t,n}$ for this joint average, the empirical envelope is defined as
$\hat{f}(\ell)=\sum_{q}\Big\langle|\mu_{t,\ell}^{(q)}|\Big\rangle_{t,n},$ (29)
@@
Empirical learnability window and empirical matched statistic.
@@
where the right-hand side is a positive constant depending only on $\alpha$ and the error probability $\epsilon$. The absolute value is appropriate because the detection problem compares the two location shifts $\pm\tfrac{1}{2}\Delta(\ell)$ in Eq. (18), so the KL and Fano bounds depend only on the unsigned separation, equivalently on $\Delta(\ell)^{2}$. The stochastic fluctuation scale of the averaged matched statistic is $\sigma_{\alpha}(\ell)/N^{1-1/\alpha}$, so $N^{1-1/\alpha}/\sigma_{\alpha}(\ell)$ is the corresponding inverse noise scale. This reformulation makes explicit that the Fano bound constrains a single scalar quantity, namely the ratio of the unsigned signal amplitude $|\Delta(\ell)|$ to that lag-dependent fluctuation scale, and that the (aggregated) detection threshold $\varepsilon$ affects only the absolute position of $\mathcal{H}_{N}$, not how it scales with $N$.
In practice, however, the population signs $\operatorname{sgn}(m_{q}(\ell))$ in (13) are not available, so we use the raw signed empirical matched statistic
$\widetilde{S}_{t,\ell}\;=\;\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\zeta^{(q)}_{t,\ell}$. (31)
Applying the same two-stage aggregation used for the theoretical construction in (16)–(17) to the raw signed statistic yields the empirical counterpart
$\widetilde{S}_{N}(\ell)\;=\;\frac{1}{N(T-\ell)}\sum_{n=1}^{N}\sum_{t=\ell+1}^{T}\widetilde{S}^{(n)}_{t,\ell}$. (32)
Because the oracle sign factors in (13) are unavailable, $\widetilde{S}_{N}(\ell)$ may have either sign even when a lag-$\ell$ signal is present. This sign ambiguity is immaterial for the detection problem in Sec. 5.4.2, where the two hypotheses differ only by opposite location shifts $\theta_{\mathrm{out}}=\pm\tfrac{1}{2}\Delta(\ell)$ in Eq. (18), so detectability depends on the magnitude of the mean shift rather than on its orientation. We therefore define the empirical unsigned signal estimate $\widehat{\Delta}(\ell)=\bigl|\widetilde{S}_{N}(\ell)\bigr|$ and the empirical signal-to-noise ratio
$\widehat{\mathrm{SNR}}(\ell,N)\;=\;\frac{\widehat{\Delta}(\ell)\;N^{1-1/\hat{\alpha}}}{\hat{\sigma}_{\alpha}(\ell)}.$ (33)
The sign-oriented theoretical matched statistic (13) removes cancellation and therefore defines an oracle benchmark; the empirical raw statistic (31) may yield a smaller signal, making the resulting learnability estimates conservative relative to that benchmark.
Since $c_{\alpha}$ is an existence constant from the LAN expansion (Appendix F) and is not available in closed form, we set the detection threshold $\varepsilon$ directly. This preserves the full scaling structure of the theoretical bound (the concentration rate $N^{1-1/\alpha}$, the dependence on signal amplitude and noise scale) while replacing the single unknown constant with a user-specified parameter.
The empirical learnability window is then
@@
For interpretability, we report the threshold through its inverse, the noise tolerance $1/\varepsilon$, which represents the maximum tolerated inverse SNR in the empirical detection criterion. Thus, $1/\varepsilon=0.1$ means detection requires $\widehat{\mathrm{SNR}}(\ell,N)>10$, while $1/\varepsilon=0.05$ requires $\widehat{\mathrm{SNR}}(\ell,N)>20$. Smaller $1/\varepsilon$ therefore imposes a stricter detectability requirement. Changing $1/\varepsilon$ adjusts the operating point of the empirical learnability window, while preserving the same scaling form in $N$, signal amplitude, and noise scale. All experiments use $1/\varepsilon=0.05$ ($\varepsilon=20$).
@@ 7 Discussion and future directions
This work develops a theoretical framework linking gating geometry, optimizer adaptation, heavy-tailed statistics, and the finite-horizon learnability of recurrent neural networks. The central message is that the generalized effective learning rates, rather than Jacobian stability in isolation, determine how much gradient signal remains statistically detectable across temporal lags. The decay profile of the envelope, together with the tail index $\alpha$, governs both sample-complexity requirements and the learnability window.
@@
The scaling laws make precise the statistical cost of each decay regime: exponential envelope decay entails exponential sample complexity, while polynomial decay entails only polynomial sample complexity. For any finite training budget, exponential forgetting rapidly exhausts the available statistical resources, whereas polynomial forgetting keeps long-range dependencies within reach. Heavy-tailed fluctuations exacerbate this disparity by increasing $\kappa_{\alpha}$ and slowing statistical concentration. Heavy-tailed noise therefore acts as a viability constraint, disfavoring rapid forgetting and effectively selecting, among the dynamical regimes accessible to an architecture–optimizer pair, those that preserve signal over longer temporal scales.
We conjecture that this viability constraint induces a form of constrained self-organization: during training, recurrent networks—and potentially deep learning models more broadly—tend to evolve toward regimes of slower envelope decay in order to maintain learnability under their own gradient-noise statistics. This does not require explicit regularization; it follows from the fact that configurations with exponential forgetting become statistically unlearnable at long lags, creating implicit selection pressure toward slower decay. The achievable regime depends on the architecture–optimizer pair: constrained gating limits the accessible spectra, while flexible gating combined with adaptive optimization permits broader time-scale mixtures. This conjecture is consistent with our empirical observation that slow-decay regimes systematically co-occur with heavy-tailed statistics, whereas the converse does not strictly hold: heavy-tailed fluctuations can coexist with concentrated or even collapsed spectra when the architecture lacks the parametric flexibility to translate stochastic forcing into spectral broadening. The robust directional pattern is that sustained slow decay requires heavy-tailed noise, but heavy-tailed noise alone is not sufficient.
@@
These observations highlight three broader implications. First, the temporal horizon of learnability is primarily controlled by the joint action of the envelope decay and optimizer adaptation. Second, the architecture–optimizer pair shapes the space of attainable time-scale spectra, biasing—but not guaranteeing—the emergence of heterogeneous dynamics that support long-range credit assignment. Third, under heavy-tailed fluctuations, increasing the number of training sequences yields diminishing returns for configurations operating in the exponential regime, whereas modifying the architecture–optimizer pair can alter the scaling class of temporal learnability entirely.